OpenStudy (s3a):

(a) What is the equation of the tangent plane to the surface z = 2xy + 3y^2 + 2x + y + 4 at the point P(-4,-3,60)? (b) What are the parametric equations of the normal line at this point?

OpenStudy (s3a):

For (a), I get z = -12x -25y - 63 For (b), I get that the constant are -4, -3, and 60 for x, y, and z respectively but, I don't know how to find the coefficients of t for x, y, and z and would appreciate it if someone could help me figure this out.

OpenStudy (s3a):

Basically, I get it all except the coefficients of t for part (b).

OpenStudy (zarkon):

you might want to try part (a) again

OpenStudy (s3a):

My answer is correct. I checked.

OpenStudy (zarkon):

did you type the question correctly....(-4,-3,60) is not a solution to the equation you gave above

OpenStudy (s3a):

Sorry, z != 2xy + 3y^2 + 2x + y + 4 z = x^2 + 2xy + 3y^2 + 2x + y + 4

OpenStudy (s3a):

(I'm sleepy.)

OpenStudy (zarkon):

do you know how to find the normal vector to the plane?

OpenStudy (s3a):

Is it -12 i - 25 j - 63k ?

OpenStudy (s3a):

and do you mean *a* normal vector?

OpenStudy (zarkon):

no that is not a normal vector to the surface

OpenStudy (s3a):

So, I don't know. :(

OpenStudy (zarkon):

\[a(x-x_0)+b(y-y_0)+c(z-z_0)=0\] is a plane with normal vector \(\vec{n}=<a,b,c>\)

OpenStudy (zarkon):

find the coefficients infront of the x,y,z variables

OpenStudy (s3a):

Oh wait! Is the normal vector: 12 i + 25 j + k ?

OpenStudy (s3a):

(I did that based on what you just said.)

OpenStudy (zarkon):

yes

OpenStudy (s3a):

Oh so the coefficients of t are the coefficients of i, j, k.

OpenStudy (s3a):

I get it then :) Thanks!

OpenStudy (zarkon):

don't forget to include your original point (-4,-3,60)